# On the shortest addition chains of numbers of special forms and progress on the Scholz conjecture

### v10

@@ Abstract.
In this paper we study the shortest addition chains of numbers of special forms. We obtain the crude inequality
$\iota(2^{n}-1)\leq n+1+G(n)$
for some function $G:\mathbb{N}\longrightarrow\mathbb{N}$. In particular we obtain the weaker inequality
$\iota(2^{n}-1)\leq n+1+\left\lfloor\frac{n-2}{2}\right\rfloor$
where $\iota(n)$ is the length of the shortest addition chain producing $n$. Additionally, we prove an inequality relating the length of addition chains producing number of the form $12^{n}-$ 1 to the length of their shortest addition chain producing their exponents. In particular, we obtain the inequality
$\delta(2^{n}-1)\leq n-1+\iota(n)+G(n)$
where $\delta(n)$ and $\iota(n)$ denotes the length of an addition chain and the shortest addition chain producing $n$, respectively, with $K:\mathbb{N}\longrightarrow\mathbb{R}$. We also obtain a more explicit improved upper bound in the form
where $\xi:\mathbb{N}\longrightarrow\mathbb{R}$. As a consequence, we obtain the inequality
$\iota(2^{n}-1)\lesssim n+\iota(n)+\frac{n}{\log n}+1.3\log n\int\limits_{2}^{\frac{n-1}{2}}\frac{dt}{\log^{3}t}+\xi(n).$
By using a slightly improved iteration method and the ”pothole” method on the factors of numbers of the form $12^{n}-$ 1, we prove the stronger inequality
for all $n\in\mathbb{N}$ with $n\geq 64$ for $10\leq\xi(n,j)<$ 1, where $\iota(\cdot)$ denotes the length of the shortest addition chain producing $\cdot$. This inequality is stronger than
$\iota(r)<\frac{\log r}{\log 2}(1+\frac{1}{\log\log r}+\frac{2\log 2}{(\log r)^{1-\log 2}})$
in the case $r=2^{n}-1$ but slightly weaker than the conjectured inequality
$\iota(2^{n}-1)\leq n-1+\iota(n).$
@@ 3. Addition chains of numbers of special forms
Corollary 3.1 could be considered as a weaker and a crude form of the inequality relating the length of the shortest addition producing $12^{n}-$ 1 to the length of the shortest addition chain producing $n$. In particular, the scholz conjecture is the claim that
@@
4. Scholz conjecture
In this section we study scholz conjecture. We prove some an inequality related to the scholz conjecture on additions chains. We begin with the following fundamental result which can be found in LABEL:brauer1939addition.
@@ 4. Scholz conjecture
We now obtain an inequality related to scholz conjecture.
@@
with $a_{i}=2^{i-2}=r_{i}$ for $12\leq i\leq n+$ 1, where $a_{i}$ and $r_{i}$ denotes the determiner and the regulator of the $i$ th generator of the chain. Let us consider only the complete sub-addition chain
@@
This result would follow from the inequality established if we can take $\delta(2^{n}-1)=\iota(2^{n}-1)$ and choose $K(n)=0$ in Theorem 4.3. In other words, the scholz conjecture can be reformulated in the following way:
@@ 4.1. Explicit upper bound
In this section, we prove an explicit upper bound for the length of the shortest addition chain producing numbers of the form $12^{n}-$ 1.
@@
with $a_{i}=2^{i-2}=r_{i}$ for $12\leq i\leq n+$ 1, where $a_{i}$ and $r_{i}$ denotes the determiner and the regulator of the $i$ th generator of the chain. Let us consider only the complete sub-addition chain
@@ 5. The backtracking method
6. The filling the Pothole method
@@ 4. $\bullet$
with $a_{i}=2^{i-2}=r_{i}$ for $12\leq i\leq n+$ 1, where $a_{i}$ and $r_{i}$ denotes the determiner and the regulator of the $i$ th generator of the chain. Let us consider only the complete sub-addition chain
@@ 6.1. Improved pothole method
terms to the original complete sub-addition chain, where the upper bound follows by virtue of Lemma LABEL:key. Since the inequality holds
@@
Now we construct an addition chain producing $12^{k}+$ 1. We construct the addition chain producing $2^{k}$ as $1,2,2^{2},\ldots,2^{k-1},2^{k}$ with corresponding sequence of partition
@@ 9. Progress on the Scholz conjecture
holds for all $n\in\mathbb{N}$ with $n\geq 64$ for $10\leq\xi(n,j)<$ 1, where $\iota(\cdot)$ denotes the length of the shortest addition chain producing normal-$\cdot$.
@@
terms to the original complete sub-addition chain, where the upper bound follows by virtue of Lemma LABEL:key. Since the inequality holds
@@ 10. The notion of carries
11. Improved inequality using the notion of carries
@@
holds for all $n\in\mathbb{N}$ with $n\geq 4$, where $\iota(\cdot)$ denotes the length of the shortest addition chain producing normal-$\cdot$.
@@
(11.7) $\iota(2^{n}-1)$ $\leq\sum\limits_{v=1}^{s}\frac{n}{2^{v}}+s+2\sum\limits_{j=1}^{s}\sum\limits_{\begin{subarray}{c}\eta(2^{m}-1)\neq 0\\ m=\lfloor\frac{n}{2^{j-1}}\rfloor\end{subarray}}1-\theta(n,s)+\iota(2^{\frac{n}{2^{s}}-\xi(n,s)}-1)$
@@
terms to the original complete sub-addition chain, where the upper bound follows by virtue of Lemma LABEL:key. Since the inequality holds
@@
(11.18) $\iota(2^{n}-1)$ $\leq\sum\limits_{v=1}^{s}\frac{n}{2^{v}}+s+2\sum\limits_{j=1}^{s}\sum\limits_{\begin{subarray}{c}\eta(2^{m}-1)\neq 0\\ m=\lfloor\frac{n}{2^{j-1}}\rfloor\end{subarray}}1-\theta(n,s)+\iota(2^{\frac{n}{2^{s}}-\xi(n,s)}-1)$
@@
(11.24) $=n-1+\frac{n}{2^{s}}+s+2\sum\limits_{j=1}^{s}\sum\limits_{\begin{subarray}{c}\eta(2^{m}-1)\neq 0\\ m=\lfloor\frac{n}{2^{j-1}}\rfloor\end{subarray}}1-\theta(n,s)-\left\lfloor\frac{\frac{n}{2^{s}}-\xi(n,s)-1}{2^{\lfloor\frac{\log n}{\log 2}\rfloor+1-s}}\right\rfloor$
@@
holds for all $n\in\mathbb{N}$ with $n\geq 4$, where $\iota(\cdot)$ denotes the length of the shortest addition chain producing normal-$\cdot$.
@@
(11.31) $\iota(2^{n}-1)$ $\leq\sum\limits_{v=1}^{s}\frac{n}{2^{v}}+s+2\sum\limits_{j=1}^{s}\sum\limits_{\begin{subarray}{c}\eta(2^{m}-1)\neq 0\\ m=\lfloor\frac{n}{2^{j-1}}\rfloor\end{subarray}}1-\theta(n,s)+\iota(2^{\frac{n}{2^{s}}-\xi(n,s)}-1)$
@@
terms to the original complete sub-addition chain, where the upper bound follows by virtue of Lemma LABEL:key. Since the inequality holds
@@
(11.42) $\iota(2^{n}-1)$ $\leq\sum\limits_{v=1}^{s}\frac{n}{2^{v}}+s+2\sum\limits_{j=1}^{s}\sum\limits_{\begin{subarray}{c}\eta(2^{m}-1)\neq 0\\ m=\lfloor\frac{n}{2^{j-1}}\rfloor\end{subarray}}1-\theta(n,s)+\iota(2^{\frac{n}{2^{s}}-\xi(n,s)}-1)$
@@
(11.48) $=n-1+\frac{n}{2^{s}}+s+2\sum\limits_{j=1}^{s}\sum\limits_{\begin{subarray}{c}\eta(2^{m}-1)\neq 0\\ m=\lfloor\frac{n}{2^{j-1}}\rfloor\end{subarray}}1-\theta(n,s)-\left\lfloor\frac{\frac{n}{2^{s}}-\xi(n,s)-1}{2^{\lfloor\frac{\log n}{\log 2}\rfloor+1-s}}\right\rfloor$
@@
for all $n\in\mathbb{N}$ with $n\geq 4$ for $10\leq\xi(n,j)<$ 1, where $\iota(\cdot)$ denotes the length of the shortest addition chain producing normal-$\cdot$.
@@
holds for all $n\in\mathbb{N}$ with $n\geq 4$, where $\iota(\cdot)$ denotes the length of the shortest addition chain producing normal-$\cdot$.
@@
holds for all $n\in\mathbb{N}$ with $n\geq 4$, where $\iota(\cdot)$ denotes the length of the shortest addition chain producing normal-$\cdot$.
The following chain of results we have obtained illustrates that to make progress on the Scholz conjecture, it suffices to study possible way of controlling numbers of the form $12^{n}-$ 1 with high carries. In other words, the degree of carries of numbers of the form $12^{n}-$ 1 determines the quality of the upper bound for its corresponding length of the shortest addition using the current method.

### v11

# On addition chains and progress on the Scholz conjecture

@@ Abstract.
In this paper, we develop some new classes of methods to study the Scholz conjecture on addition chains. It turns out that the exponents of numbers of the form $12^{n}-$ 1 largely determine the length of the shortest addition chain for number producing $12^{n}-$ 1. Exploiting the notion of carries, we obtain improved upper bounds for the length of the shortest addition chains $\iota(2^{n}-1)$ producing $12^{n}-$ 1. Most notably, we show that if $12^{n}-$ 1 has carries of degree at most
$\kappa(2^{n}-1)=\frac{1}{2}(\iota(n)-\lfloor\frac{\log n}{\log 2}\rfloor+\sum\limits_{j=1}^{\lfloor\frac{\log n}{\log 2}\rfloor}\{\frac{n}{2^{j}}\})$
then the inequality
$\iota(2^{n}-1)\leq n+1+\sum\limits_{j=1}^{\lfloor\frac{\log n}{\log 2}\rfloor}\bigg{(}\{\frac{n}{2^{j}}\}-\xi(n,j)\bigg{)}+\iota(n)$
[…]
@@ 3. Addition chains of numbers of special forms
Corollary 3.1 could be considered as a weaker and a crude form of the inequality relating the length of the shortest addition producing $12^{n}-$ 1 to the length of the shortest addition chain producing $n$. In particular, Scholz’s conjecture is the claim that
@@ 3.1. The symmetry method
4. Length of general addition chains
In this section we study Scholz’s conjecture. We prove some an inequality related to the conjecture on additions chains. We begin with the following fundamental result which can be found in [1].
@@ 4. Length of general addition chains
We now obtain an inequality related to Scholz’s conjecture.
@@
with $a_{i}=2^{i-2}=r_{i}$ for $12\leq i\leq n+$ 1, where $a_{i}$ and $r_{i}$ denotes the determiner and the regulator of the $i^{th}$ generator of the chain. Let us consider only the complete sub-addition chain
@@
This result would follow from the inequality established if we can take $\delta(2^{n}-1)=\iota(2^{n}-1)$ and choose $K(n)=0$ in Theorem 4.3. In other words, Scholz’s conjecture can be reformulated in the following way:
@@ 4.1. Explicit upper bound
In this section, we prove an explicit upper bound for the length of an addition chain - not necessarily the shortest - producing numbers of the form $12^{n}-$ 1. In particular, we study the problem of constructing an addition chain whose length meets a specified threshold requirement.
@@
with $a_{i}=2^{i-2}=r_{i}$ for $12\leq i\leq n+$ 1, where $a_{i}$ and $r_{i}$ denotes the determiner and the regulator of the $i^{th}$ generator of the chain. Let us consider only the complete sub-addition chain
@@ 5. The backtracking method
6. Filling the Pothole method
@@ 4. $\bullet$
with $a_{i}=2^{i-2}=r_{i}$ for $12\leq i\leq n+$ 1, where $a_{i}$ and $r_{i}$ denotes the determiner and the regulator of the $i^{th}$ generator of the chain. Let us consider only the complete sub-addition chain
@@ 6.1. Improved pothole method
terms to the original complete sub-addition chain, where the upper bound follows by virtue of Lemma 4.1. Since the inequality holds
@@
Now, we construct an addition chain producing $12^{k}+$ 1. We construct the addition chain producing $2^{k}$ as $1,2,2^{2},\ldots,2^{k-1},2^{k}$ with corresponding sequence of partition
@@ 9. Progress on the Scholz conjecture
[…]
@@
and so on. That is, $\xi(n,1):=\{\frac{n}{2}\}$ with $\xi(n,2)=\{\frac{1}{2}\lfloor\frac{n}{2}\rfloor\}$ and so on. Indeed the function $\xi(n,s)$ for values of $s\geq 3$ can be read from exponents of the terms arising from the iteration process. It follows from (11.26) the inequality
@@
terms to the original complete sub-addition chain. Since the inequality holds
@@ 10. The notion of carries
11. Improved inequality using the method of carries
@@
holds for all $n\in\mathbb{N}$ with $n\geq 4$, where $\iota(\cdot)$ denotes the length of the shortest addition chain producing $\cdot$.
@@
and so on. That is, $\xi(n,1):=\{\frac{n}{2}\}$ with $\xi(n,2)=\{\frac{1}{2}\lfloor\frac{n}{2}\rfloor\}$ and so on. Indeed the function $\xi(n,s)$ for values of $s\geq 3$ can be read from exponents of the terms arising from the iteration process. It follows from (11.26) the inequality
$\iota(2^{n}-1)$ $\leq\sum\limits_{v=1}^{s}\frac{n}{2^{v}}+s+2\sum\limits_{j=1}^{s}\sum\limits_{\begin{subarray}{c}\eta(2^{m}-1)\neq 0\\ m=\lfloor\frac{n}{2^{j-1}}\rfloor\end{subarray}}1-\theta(n,s)+\iota(2^{\frac{n}{2^{s}}-\xi(n,s)}-1)$
@@
terms to the original complete sub-addition chain. Since the inequality holds
@@
$\iota(2^{n}-1)$ $\leq\sum\limits_{v=1}^{s}\frac{n}{2^{v}}+s+2\sum\limits_{j=1}^{s}\sum\limits_{\begin{subarray}{c}\eta(2^{m}-1)\neq 0\\ m=\lfloor\frac{n}{2^{j-1}}\rfloor\end{subarray}}1-\theta(n,s)+\iota(2^{\frac{n}{2^{s}}-\xi(n,s)}-1)$
@@
$=n-1+\frac{n}{2^{s}}+s+2\sum\limits_{j=1}^{s}\sum\limits_{\begin{subarray}{c}\eta(2^{m}-1)\neq 0\\ m=\lfloor\frac{n}{2^{j-1}}\rfloor\end{subarray}}1-\theta(n,s)-\left\lfloor\frac{\frac{n}{2^{s}}-\xi(n,s)-1}{2^{\lfloor\frac{\log n}{\log 2}\rfloor+1-s}}\right\rfloor$
@@
holds for all $n\in\mathbb{N}$ with $n\geq 4$, where $\iota(\cdot)$ denotes the length of the shortest addition chain producing $\cdot$.
@@
and so on. That is, $\xi(n,1):=\{\frac{n}{2}\}$ with $\xi(n,2)=\{\frac{1}{2}\lfloor\frac{n}{2}\rfloor\}$ and so on.Indeed the function $\xi(n,s)$ for values of $s\geq 3$ can be read from exponents of the terms arising from the iteration process. It follows from (11.26) the inequality
$\iota(2^{n}-1)$ $\leq\sum\limits_{v=1}^{s}\frac{n}{2^{v}}+s+2\sum\limits_{j=1}^{s}\sum\limits_{\begin{subarray}{c}\eta(2^{m}-1)\neq 0\\ m=\lfloor\frac{n}{2^{j-1}}\rfloor\end{subarray}}1-\theta(n,s)+\iota(2^{\frac{n}{2^{s}}-\xi(n,s)}-1)$
@@
terms to the original complete sub-addition chain. Since the inequality holds
@@
$\iota(2^{n}-1)$ $\leq\sum\limits_{v=1}^{s}\frac{n}{2^{v}}+s+2\sum\limits_{j=1}^{s}\sum\limits_{\begin{subarray}{c}\eta(2^{m}-1)\neq 0\\ m=\lfloor\frac{n}{2^{j-1}}\rfloor\end{subarray}}1-\theta(n,s)+\iota(2^{\frac{n}{2^{s}}-\xi(n,s)}-1)$
@@
$=n-1+\frac{n}{2^{s}}+s+2\sum\limits_{j=1}^{s}\sum\limits_{\begin{subarray}{c}\eta(2^{m}-1)\neq 0\\ m=\lfloor\frac{n}{2^{j-1}}\rfloor\end{subarray}}1-\theta(n,s)-\left\lfloor\frac{\frac{n}{2^{s}}-\xi(n,s)-1}{2^{\lfloor\frac{\log n}{\log 2}\rfloor+1-s}}\right\rfloor$
@@
for all $n\in\mathbb{N}$ with $n\geq 4$ for $10\leq\xi(n,j)<$ 1, where $\iota(\cdot)$ denotes the length of the shortest addition chain producing $\cdot$.
@@
holds for all $n\in\mathbb{N}$ with $n\geq 4$, where $\iota(\cdot)$ denotes the length of the shortest addition chain producing $\cdot$.
@@
holds for all $n\in\mathbb{N}$ with $n\geq 4$, where $\iota(\cdot)$ denotes the length of the shortest addition chain producing $\cdot$.
The following chain of results we have obtained illustrates that to make progress on the Scholz conjecture, it suffices to study possible way of controlling numbers of the form $12^{n}-$ 1 with high carries. In other words, the degree of carries of numbers of the form $12^{n}-$ 1 determines the quality of the upper bound for its corresponding length of the shortest addition using the current method. We end the paper by proving using the same method that with the expected number of carries for a fixed number $12^{n}-$ 1, we can obtain the stronger result:
Theorem 11.6.
If $12^{n}-$ 1 has carries of degree at most
$\kappa(2^{n}-1)=\frac{1}{2}(\iota(n)-\lfloor\frac{\log n}{\log 2}\rfloor+\sum\limits_{j=1}^{\lfloor\frac{\log n}{\log 2}\rfloor}\{\frac{n}{2^{j}}\})$
then the inequality
$\iota(2^{n}-1)\leq n+1+\sum\limits_{j=1}^{\lfloor\frac{\log n}{\log 2}\rfloor}\bigg{(}\{\frac{n}{2^{j}}\}-\xi(n,j)\bigg{)}+\iota(n)$
[…]
Proof.
Suppose that $12^{n}-$ 1 has at most
$\frac{1}{2}(\iota(n)-\lfloor\frac{\log n}{\log 2}\rfloor+\sum\limits_{j=1}^{\lfloor\frac{\log n}{\log 2}\rfloor}\{\frac{n}{2^{j}}\})$
where
$\eta(2^{n}-1):=\frac{(1-(-1)^{n})}{2}(2^{n-(1-(-1)^{n})\frac{1}{2}})$
$2^{n}-1=(2^{\frac{n}{2}}-1)(2^{\frac{n}{2}}+1)$
and
$2^{n}-1=(2^{\frac{n-1}{2}}-1)(2^{\frac{n-1}{2}}+1)+2^{n-1}$
if $n\equiv 1\pmod{2}$. By combining both cases, we obtain the inequality
Applying Lemma 6.3, we obtain further the inequality
where
$\eta(2^{k}-1)=\frac{(1-(-1)^{k})}{2}(2^{k-(1-(-1)^{k})\frac{1}{2}})$
$2^{k}-1=(2^{\frac{k}{2}}-1)(2^{\frac{k}{2}}+1)$
and
$2^{k}-1=(2^{\frac{k-1}{2}}-1)(2^{\frac{k-1}{2}}+1)+2^{k-1}$
if $k\equiv 1\pmod{2}$. By combining both cases, we obtain the inequality
Applying Lemma 6.3, we obtain further the inequality
so that by inserting (11.24) into (11.23), we obtain the inequality
(11.25) $+\iota(2^{\lfloor\frac{n}{2}\rfloor}+1)+\eta(2^{n}-1).$
Next, we iterate the factorization up to frequency $s$ to obtain
$\xi(n,1)=(1-(-1)^{n})\frac{1}{4}<1$
and
$\xi(n,2)=(1-(-1)^{n})\frac{1}{8}+(1-(-1)^{k})\frac{1}{4}<1$
with
$k:=\lfloor\frac{n}{2}\rfloor$
$\iota(2^{n}-1)$ $\leq\sum\limits_{v=1}^{s}\frac{n}{2^{v}}+s+2\sum\limits_{j=1}^{s}\sum\limits_{\begin{subarray}{c}\eta(2^{m}-1)\neq 0\\ m=\lfloor\frac{n}{2^{j-1}}\rfloor\end{subarray}}1-\theta(n,s)+\iota(2^{\frac{n}{2^{s}}-\xi(n,s)}-1)$
where the term
$\theta(n,s):=\sum\limits_{j=1}^{s}\xi(n,j)=0$
$2=1+1,2+2=2^{2},2^{2}+2^{2}=2^{3},\ldots,2^{k-1}=2^{k-2}+2^{k-2}.$
terms to the original complete sub-addition chain. Since the inequality holds
$=2^{k}-2$
we insert terms into the sum
so that we have
$\sum\limits_{i=1}^{k-1}2^{i}=2^{k}-2.$
$k-2-\lfloor\frac{k-1}{2}\rfloor$
$\lfloor\frac{k-1}{2}\rfloor-\lfloor\frac{k-1}{2^{2}}\rfloor-1$
(11.30) $\cdots+2^{1}.$
$\lfloor\frac{k-1}{2^{j}}\rfloor-\lfloor\frac{k-1}{2^{j+1}}\rfloor-1$
obtained by adding the numbers in the chain
$k-1-\lfloor\frac{k-1}{2}\rfloor-1$
$\lfloor\frac{k-1}{2}\rfloor-\lfloor\frac{k-1}{2^{2}}\rfloor-1$
[…]
$\iota(2^{n}-1)$ $\leq\sum\limits_{v=1}^{s}\frac{n}{2^{v}}+s+2\sum\limits_{j=1}^{s}\sum\limits_{\begin{subarray}{c}\eta(2^{m}-1)\neq 0\\ m=\lfloor\frac{n}{2^{j-1}}\rfloor\end{subarray}}1-\theta(n,s)+\iota(2^{\frac{n}{2^{s}}-\xi(n,s)}-1)$
$=n-1+\frac{n}{2^{s}}+s+2\sum\limits_{j=1}^{s}\sum\limits_{\begin{subarray}{c}\eta(2^{m}-1)\neq 0\\ m=\lfloor\frac{n}{2^{j-1}}\rfloor\end{subarray}}1-\theta(n,s)-\left\lfloor\frac{\frac{n}{2^{s}}-\xi(n,s)-1}{2^{\lfloor\frac{\log n}{\log 2}\rfloor+1-s}}\right\rfloor$
where we note that
and we obtained that
$\sum\limits_{j=1}^{s}\sum\limits_{\begin{subarray}{c}\eta(2^{m}-1)\neq 0\\ m=\lfloor\frac{n}{2^{j-1}}\rfloor\end{subarray}}1\leq\frac{1}{2}(\iota(n)-\lfloor\frac{\log n}{\log 2}\rfloor+\sum\limits_{j=1}^{\lfloor\frac{\log n}{\log 2}\rfloor}\{\frac{n}{2^{j}}\})$
and the inequality
$\iota(2^{n}-1)\leq n-1-\theta(n,\lfloor\frac{\log n}{\log 2}\rfloor)+\lfloor\frac{\log n}{\log 2}\rfloor+2+\iota(n)-\lfloor\frac{\log n}{\log 2}\rfloor+\sum\limits_{j=1}^{\lfloor\frac{\log n}{\log 2}\rfloor}\{\frac{n}{2^{j}}\}$
[…]
It turns out that proving integers of the form $12^{n}-$ 1 has carries of degree at most
$\kappa(2^{n}-1)=\frac{1}{2}(\iota(n)-\lfloor\frac{\log n}{\log 2}\rfloor+\sum\limits_{j=1}^{\lfloor\frac{\log n}{\log 2}\rfloor}\{\frac{n}{2^{j}}\})$
would yield the inequality
$\iota(2^{n}-1)\leq n+1+\sum\limits_{j=1}^{\lfloor\frac{\log n}{\log 2}\rfloor}\bigg{(}\{\frac{n}{2^{j}}\}-\xi(n,j)\bigg{)}+\iota(n)$
for all $n\in\mathbb{N}$ as shown in the preceding proof, which is slightly short of the original conjecture. Indeed, we expect the degree of carries of all integers to be of the above form since for integers of the form $n=2^{k}$, it matches exactly with degree given by the formula. In particular the number $12^{2^{k}}-$ 1 is always free of carries and we see that
$\kappa(2^{2^{k}}-1)=\frac{1}{2}(\iota(2^{k})-\lfloor\frac{\log 2^{k}}{\log 2}\rfloor+\sum\limits_{j=1}^{\lfloor\frac{\log n}{\log 2}\rfloor}\{\frac{2^{k}}{2^{j}}\})=0.$
We make the following conjecture
Conjecture 11.1.
(The carry-addition chain conjecture) Let $n\geq 2$ and let $\kappa(\cdot)$ denotes the degree of carry of $\cdot$. Then we have
$\kappa(2^{n}-1)=\bigg{\lceil}\frac{1}{2}(\iota(n)-\lfloor\frac{\log n}{\log 2}\rfloor+\sum\limits_{j=1}^{\lfloor\frac{\log n}{\log 2}\rfloor}\{\frac{n}{2^{j}}\})\bigg{\rceil}$
[…]